OpenStudy (anonymous):

What is the x-coordinate of the x-intercept of the line that passes through points (-5, 3) and (1, -3)?

OpenStudy (anonymous):

you find the line for those points....graph it...and then look at the x-intercept! :)

OpenStudy (anonymous):

can ya help me or someone please

OpenStudy (anonymous):

so the slope is (-3-3)/(1+5) which is?

OpenStudy (anonymous):

solve (-3-3)/(1+5) and you get -6/6 so simplify that and you get...?

OpenStudy (anonymous):

-2,0

OpenStudy (anonymous):

no no....what is -6/6? :)

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

so we have a slope (-1) now lets find the equation for the line to do that we plug in the slope(m) into the equation y=mx+b and we put in any one of the two given values of x an y :) so like: (3)=(-1)(-5)+b

OpenStudy (anonymous):

so wht now

OpenStudy (anonymous):

solve that for b and you get b=-2 so you use that for your equation: y=-x-2 now you graph that and look for the x-intercept! :)

OpenStudy (anonymous):

wait huh

OpenStudy (anonymous):

so y = -2?

OpenStudy (anonymous):

y = m x + b y = -1 x + (-2)

OpenStudy (anonymous):

so now i graph that

OpenStudy (anonymous):

-2 is x

OpenStudy (anonymous):

heres how to graph it: